OpenStudy (anonymous):

divide 63x/7x5 A. 9/x4 B. 9/x5 C. 1/9x4 D. 9/x4

OpenStudy (anonymous):

A. 9/x4 is the correct answer. @LaBlueEyes2013

OpenStudy (tkhunny):

You have written: \(\dfrac{63x}{7}\cdot x^{5}\). Did you mean \(\dfrac{63x}{7x^{5}}\)?

OpenStudy (anonymous):

the second one

OpenStudy (tkhunny):

Please remember your Order of Operations and add parentheses where necessary. \(\dfrac{63x}{7x^{5}} = \dfrac{63}{7}\cdot\dfrac{x}{x^{5}} = 9\cdot x^{1-5}\) What else?